OpenStudy (anonymous):

How do I rewrite this in standard form? 4x^2+24x+128y-988=0

OpenStudy (jingle535):

Standard form of a quadratic equation is ax² + bx + c = 0 Since that has a y I am not sure it can be put in standard form unless you meant for that to be an x, in which case it would be 4x2+152x−988=0.

OpenStudy (anonymous):

Sorry, I meant to specify standard form for a parabola.

OpenStudy (anonymous):

standard form of parabola's look like y = ax^2 + bx + c so get y to one side

OpenStudy (jingle535):

I was just typing that out lol :)

OpenStudy (jingle535):

I forgot about parabolas sorry fkadan ;)

OpenStudy (anonymous):

It's ok, thanks for the help!!